OpenStudy (anonymous):

Multiply this rational expression, 7x/14x+7 * 8x+4/5

OpenStudy (anonymous):

Could you please add some brackets to the expression so as to show it clearly?

OpenStudy (anonymous):

sorry (7x)/(14x+7) * (8x+4)/(5)

OpenStudy (anonymous):

this is the product of two fraction?

OpenStudy (anonymous):

yes I think so :)

OpenStudy (anonymous):

(7x)/(14x+7) * (8x+4)/(5) =x/(2x+1) * 4(2x+1)/5 =4x/5

OpenStudy (anonymous):

You're correct thank you !!

OpenStudy (anonymous):

you are welcome! :D